# Aligning Constraint Generation with Design Intent in Parametric CAD

### v1

@@ Abstract
We adapt alignment techniques from reasoning LLMs to the task of generating engineering sketch constraints found in computer-aided design (CAD) models. Engineering sketches consist of geometric primitives (e.g. points, lines) connected by constraints (e.g. perpendicular, tangent) that define the relationships between them. For a design to be easily editable, the constraints must effectively capture design intent, ensuring the geometry updates predictably when parameters change. Although current approaches can generate CAD designs, an open challenge remains to align model outputs with design intent, we label this problem ‘design alignment’. A critical first step towards aligning generative CAD models is to generate constraints which fully-constrain all geometric primitives, without over-constraining or distorting sketch geometry. Using alignment techniques to train an existing constraint generation model with feedback from a constraint solver, we are able to fully-constrain 93% of sketches compared to 34% when using a naïve supervised fine-tuning (SFT) baseline and only 8.9% without alignment. Our approach can be applied to any existing constraint generation model and sets the stage for further research bridging alignment strategies between the language and design domains. Additional results can be found in the supplementary materials.
[…]
@@ 2 Related Work
Design Alignment  Beyond the language domain, alignment techniques have been used to improve and align image generation. Lee et al. [18] propose fine-tuning diffusion-based text-to-image models using human feedback, significantly improving alignment between textual prompts and generated visuals. Similarly, ImageReward [40] uses a learned reward model trained on human preference data, which guides the diffusion model fine-tuning toward images preferred by human evaluators. Extending this idea, Black et al. [6] reframes image generation as a sequential RL task, introducing Denoising Diffusion Policy Optimization (DDPO) to optimize complex user-defined objectives directly for alignment without explicit human annotation. Within the design domain, few works applied generative modeling alignment techniques to produce outcomes that maintain design intent. GearFormer [11] leveraged differentiable sampling to enforce preferences in the solutions for mechanical configuration design problems, however, this approach does not work with rewards that require blackbox solvers in the loop. In concurrent work, e-SimFT [9] leveraged preference data obtained from physics simulations to enhance the exploration of the Pareto front in a multi-preference setting, improving the optimality of solutions generated with GearFormer. We believe the problem of design alignment will become a critical part of future generative models aimed at design applications.
@@ 3 Problem
Our objective is to align the model toward generating constraint sets that yield fully-constrained sketches while minimizing cases of under-constrained, over-constrained, not solvable, or instability. This is a necessary prerequisite toward the ultimate goal of generating constraints that preserve the original sketch design intent when the designer varies the geometric parameters.
@@ 4.1 Constraint Solver
To evaluate model-generated constraints, we integrate the commercially available constraint solver in Autodesk Fusion. This provides industrial-grade accuracy and robustness, allowing us to assess sketch conditions defined in Section 3, whether applying a given set of constraints yields a stable, fully-constrained sketch or leads to over-constraint and unsolvability. Additionally, the Fusion constraint solver adjusts geometry to resolve all constraints, enabling alignment checks with design intent. On average the constraint solver takes 0.1-0.2 seconds to evaluate a sketch, although difficult sketches can be in the tens of seconds. For computational efficency we automatically deem the sketch unsolvable if the sketch takes longer than two seconds to solve.
@@ 4.3 Preference-Based Optimization (PO)
Expert Iteration (ExIt)  alternates between expert improvement and policy distillation. Following [33, 14], we use temperature sampling combined with rejection sampling to generate high-quality candidate constraint sequences. Specifically, during the exploration step, we initialize the policy model $\pi_{\theta_{t=0}}$ from the SFT model, and for each sketch query $q$ in the training set, we sample $K=8$ candidate constraint sequences $\tau$ at temperature $T=1.0$. The training dataset is constructed by discarding sequences that are under-constrained, over-constrained, or unsolvable solutions. This process is repeated $N=2$ times over the dataset, and the policy is trained using cross-entropy loss:
$\mathbb{E}_{(q,\tau)\sim\mathcal{D}}\left[\log\pi_{\theta_{t}}(\tau|q)\right]$ (1)
@@
Direct Preference Optimization (DPO)  learns from pairwise preference data, approximating an implicit reward via a reparameterized Bradley-Terry model [25]. Similar to ExIt, we construct the training dataset by sampling $K=8$ constraint sequence completions $\tau$ for each sketch query $q$ from the policy model $\pi_{\theta_{t}}$ at temperature $T=1.0$. Pairs $(\tau_{w},\tau_{l})$ are ranked based on the differences in the percentage of fully-constrained curves between $\tau_{w}$ and $\tau_{l}$. The optimization objective is:
@@ 4.4.3 Policy gradient methods
$\mathop{\mathbb{E}}_{\tau\sim\pi}\left[(r-r^{*})\nabla\log\pi_{\theta}(\tau|q)\right].\$ (3)
@@
$\mathop{\mathbb{E}}_{\{\tau\}\sim\pi}\left[\frac{1}{G}\sum_{g=1}^{G}\left[\left(r_{g}-\text{mean}(\{r_{i}\}_{i\neq g}^{G})\right)\nabla\log\pi_{\theta}(\tau_{g}|q_{g})\right]\right].\$ (4)
@@ 5.2 Quantitative Results
Figure 5: Visual comparison of sketches with constraints generated using a baseline model and with different post-training approaches applied. Curves are colored black when constrained and blue when not.
@@
Table 2 demonstrates the evaluation result of our methods in a few-shot inference setting, where the model has $K$ attempts to generate a valid solution. Since the constraint solver can verify whether a solution is correct, this scenario reflects real-world use cases where the goal is to maximize the likelihood of producing an acceptable solution within a fixed inference budget of $K$ samples (Pass@$K$). We find that while the policy gradient-based RL methods still have the overall highest performance, increasing the number of samples $K$ has a comparatively small impact on performance compared to the other methods.
@@ 5.3 Qualitative Results
We present qualitative comparisons of sketches generated by different alignment methods in Figure 5. Curves are colored black when constrained and blue when not. Methods leveraging constraint solver feedback (Iterative DPO, ExIt, ReMax, RLOO, GROO) demonstrate a promising trend towards generating fully-constrained sketches, whereas the Base and SFT models typically leave sketches under-constrained. However, across different alignment algorithms, we observe substantial variance in the degree of geometric distortion introduced by the aligned models.
@@
In summary, our results indicate that models more easily fully-constrain sketches consisting primarily of horizontal and vertical lines without distortion. In contrast, sketches involving oblique lines increase the challenge of maintaining stability, and arcs further complicate achieving fully-constrained status. Among the tested methods, RLOO and GRPO exhibit the strongest overall performance in constraint satisfaction and geometric stability. Additional qualitative results can be found in the supplementary materials.
@@ 6 Limitations
We now outline several limitations of our work suitable for future research. First, our post-training approach has not been optimized for speed. Currently it takes $\sim$3 days to train a single epoch with the SketchGraphs dataset using a 8xH100 GPU configuration. This is primarily due to the frequent interactions with the CPU-based constraint solver and the fact that solve times can be highly varied. Roughly half of the training time is spent on GPU computation and half on detokenization and solver interaction. We expect custom optimizations could significantly reduce training time. Second, our experiments focus on sketches of moderate complexity, with a maximum of 16 geometric primitives and 64 constraints. Evaluating performance on larger, more complex sketches requires further exploration. Finally, our alignment approach currently uses feedback signals from a constraint solver to represent design intent. Incorporating subjective human preferences and explicit design intent into alignment objectives remains a promising area for future work.
@@ 7 Conclusion
We demonstrated the adaption of alignment strategies from language modeling to preserve design intent in parametric CAD sketches. By using feedback from a constraint solver as a learning signal, we show the feasibility and value of alignment in parametric CAD tasks. In doing so, we pave the way for future AI-assisted design tools that incorporate design aligment.
@@ A.1 Background
A.2 Fusion Sketch Representation
The Fusion format organizes sketch elements into a hierarchical, structured representation, wherein a sketch is defined by a set of parametric geometric primitives and a set of explicit constraints between those primitives. Each geometric primitive (line, arc, circle, point, etc.) is described by its intrinsic parameters (e.g., endpoint coordinates for a line, center and radius for a circle). Alongside the primitives, the sketch includes constraints (e.g., coincident points, perpendicular or parallel lines) that impose geometric relationships to be satisfied simultaneously. These constraints serve to preserve design intent: for instance, a coincidence constraint can lock the endpoint of a line onto a circle’s circumference, or an equal-length constraint can enforce that two segments remain the same length.
@@ A.3 Sketch Tokenization
A.4 SketchGraphs Dataset
While the main body describes how the SketchGraphs dataset was filtered and converted, additional details regarding the motivation and practical considerations of each step are provided here. The primary goal of these refinements is to produce a clean, representative subset of sketches and ensure each example aligns with standard engineering constraints.
In Table A.1 we list out the supported constraint and dimension types in Onshape terminology that we included in the training data. Notably, we filter out less prevelant constraints (Symmetric, Normal, Pattern) and dimensions (CenterLine, Projected) to focus the learning task on the core geometric relationship types which form the backbone of sketch geometry. By removing these non-core constraints, we simplify the constraint vocabulary the model must learn while still covering the vast majority of design intent in sketches.
@@ Appendix B Architecture and Experiment Details
B.1 Vitruvion
@@
B.2 Preference-based Optimization Algorithms
@@ B.3 RL algorithms
KL regularization $\beta$ 0.01
Policy clipping threshold $\epsilon$ 0.2

### v2

@@ Abstract
We adapt alignment techniques from reasoning LLMs to the task of generating engineering sketch constraints found in computer-aided design (CAD) models. Engineering sketches consist of geometric primitives (e.g. points, lines) connected by constraints (e.g. perpendicular, tangent) that define the relationships between them. For a design to be easily editable, the constraints must effectively capture design intent, ensuring the geometry updates predictably when parameters change. Although current approaches can generate CAD designs, an open challenge remains to align model outputs with design intent, we label this problem ‘design alignment’. A critical first step towards aligning generative CAD models is to generate constraints which fully-constrain all geometric primitives, without over-constraining or distorting sketch geometry. Using alignment techniques to train an existing constraint generation model with feedback from a constraint solver, we are able to fully-constrain 93% of sketches compared to 34% when using a naïve supervised fine-tuning (SFT) baseline and only 8.9% without SFT. Our approach can be applied to any existing constraint generation model and sets the stage for further research bridging alignment strategies between the language and design domains. Additional results can be found at https://autodeskailab.github.io/aligning-constraint-generation/.
Figure 1: Left: a constraint solver is used to score model generated constraints $\texttt{A},\texttt{B},\texttt{C}\sim\pi_{\text{base}}$ (and $\texttt{A}’,\texttt{B}’,\texttt{C}’\sim\pi_{\text{aligned}}$). Starting with the base model $\pi_{\text{base}}$, we post-train an aligned model $\pi_{\text{aligned}}$ from this feedback. Right: Blue lines show original primitives and gray lines show geometric distortion when dimensions vary. The aligned model $\pi_{\text{aligned}}$ produces fully-constrained sketches that preserve relative geometric relationships, whereas the base model $\pi_{\text{base}}$ produces under-constrained sketches that may distort the geometry in unintended ways.
@@ 2 Related Work
Design Alignment  Beyond the language domain, alignment techniques have been used to improve and align image generation. Lee et al. [21] propose fine-tuning diffusion-based text-to-image models using human feedback, significantly improving alignment between textual prompts and generated visuals. Similarly, ImageReward [43] uses a learned reward model trained on human preference data, which guides the diffusion model fine-tuning toward images preferred by human evaluators. Extending this idea, Black et al. [6] reframes image generation as a sequential RL task, introducing Denoising Diffusion Policy Optimization (DDPO) to optimize complex user-defined objectives directly for alignment without explicit human annotation. Few works have applied alignment techniques in the design domain. GearFormer [12] used differentiable sampling to enforce preferences in the solutions for mechanical configuration design problems, however, this approach does not work with rewards that require blackbox solvers in the loop. In concurrent work, e-SimFT [10] used preference data obtained from physics simulations to enhance the exploration of the Pareto front in a multi-preference setting, improving the solutions generated with GearFormer.
@@ 3 Problem
Our objective is to align the model toward generating constraint sets that yield fully-constrained sketches while minimizing cases of under-constrained, over-constrained, not solvable, or instability. This is a necessary prerequisite toward the ultimate goal of generating constraints that preserve the original sketch design intent when the designer varies the geometric parameters. A formal definition of FC, UC, and OC is described in  [17, 16] and in the appendix.
@@ 4.1 Constraint Solver
We evaluate generated constraints using the commercially available constraint solver in Autodesk Fusion. This solver is treated as a black box, from which we retrieve sketch conditions as defined in Section 3. A general algorithmic approach commonly used in constraint solvers is described in [8]. The solver outputs the fully-constrained status for each entity if the sketch is solvable. Otherwise it classifies the sketch as over-constrained or unsolvable. The solver adjusts geometry to resolve all constraints which we compare with the original sketch to determine if the constraints caused geometry distortion (instability). On average, it takes 0.1–0.2 seconds to per solve, although complex cases may take tens of seconds. Any sketch taking longer than two seconds to solve is automatically deemed unsolvable.
@@ 4.3 Preference-Based Optimization (PO)
Expert Iteration (ExIt)  alternates between expert improvement and policy distillation. Following [36, 15], we use temperature sampling combined with rejection sampling to generate high-quality candidate constraint sequences. During the exploration step we initialize the policy model $\pi_{\theta_{t=0}}$ from the SFT model and sample $K=8$ candidate constraint sequences $\tau$ at temperature $T=1.0$ for each sketch query $q$ in the initial training set $\mathcal{D}$}. A new training dataset $\mathcal{D}^{*}_{i\in{N}}$ is constructed by discarding sequences that are under-constrained, over-constrained, or unsolvable solutions. This process is repeated $N=2$ times over the dataset, and the policy is trained using cross-entropy loss:
$\mathbb{E}_{(q,\tau)\sim\mathcal{D}^{*}_{i\in{N}}}\left[\log\pi_{\theta_{t}}(\tau|q)\right]$ (1)
@@
Direct Preference Optimization (DPO)  learns from pairwise preference data, approximating an implicit reward via a reparameterized Bradley-Terry model [28]. Similar to ExIt, we construct the training dataset by sampling $K=8$ constraint sequence completions $\tau$ for each sketch query $q$ from the policy model $\pi_{\theta_{t}}$ at temperature $T=1.0$. Pairs $(\tau_{w},\tau_{l})$ are ranked based on the differences in the percentage of fully-constrained curves between $\tau_{w}$ (preferred sequence), and $\tau_{l}$ (non-preferred sequence). The optimization objective is:
@@ 4.4.3 Policy gradient methods
$\mathop{\mathbb{E}}_{\tau\sim\pi}\left[(r-r^{*})\nabla\log\pi_{\theta}(\tau|q)\right].\\$ (3)
@@
$\mathop{\mathbb{E}}_{\{\tau\}\sim\pi}\left[\frac{1}{G}\sum_{g=1}^{G}\left[\left(r_{g}-\text{mean}(\{r_{i}\}_{i\neq g}^{G})\right)\nabla\log\pi_{\theta}(\tau_{g}|q_{g})\right]\right].\\$ (4)
@@ 5.2 Quantitative Results
Figure 5: Visual comparison of solved sketches across the aforementioned baseline and post-training methods. Curves are colored black when fully-constrained and blue when not. Overlaid red lines are the original input sketch for reference.
@@
Table 2 demonstrates the evaluation result of our methods in a few-shot inference setting, where the model has $K$ attempts to generate a fully-constrained and stable output. This scenario reflects real-world use cases where the goal is to maximize the likelihood of producing an acceptable solution within a fixed inference budget of $K$ samples (Pass@$K$). We find that while the RL-based methods still have the overall highest performance, increasing the number of samples $K$ has a comparatively small impact on performance compared to the other methods. Additional results and analysis on the impact of sampling parameters such as temperature and top-$p$ are available in the appendix.
@@ 5.3 Qualitative Results
We randomly select sketches from the test set and show the results generated by different alignment methods in  Figure 5. Curves are colored black when constrained and blue when not. Methods leveraging alignment techniques demonstrate a promising trend towards generating fully-constrained sketches, whereas the base Vitruvion and SFT models typically leave sketches under-constrained. However, across different alignment algorithms, we observe substantial variance in the degree of geometric distortion introduced by the aligned models.
@@
Our results indicate that models more easily fully-constrain sketches consisting primarily of horizontal and vertical lines without distortion. In contrast, sketches involving oblique lines increase the challenge of maintaining stability, and arcs further complicate achieving fully-constrained status. Among the tested methods, RLOO and GRPO exhibit the strongest overall performance in achieving fully-constrained status and geometric stability.
Upon closer inspection of the generated constraints for the RL-based methods, we find that these models learns to “reward hack” by adding far more dimensions than constraints. This allows the model to avoid the instability penalty because dimensions cannot cause geometry distortion. In order to counteract this behavior, we added an additional reward penalty on the ratio of constraints versus dimensions which fixes the over-dimensioning problem at the cost of reducing stability and fully-constrained status. Additional details are provided in the appendix.
To further evaluate the efficacy of each method, we conduct a forced-choice perceptual study with professional CAD designers. We show them paired images of the same sketch with different constraints applied from each method, to understand which set of constraints they prefer. We find that the preference-based optimization methods (ExIT, DPO) and RLOO with the over-dimensioning penalty consistently outperform the SFT and vanilla RLOO results. Results of this study are provided in the appendix.
@@ 6 Limitations
Our experiments focus on sketches of moderate complexity, with a maximum of 16 geometric primitives and 64 constraints. Evaluating performance on more complex sketches requires further exploration. Additionally, our alignment approach only uses feedback signals from a constraint solver to represent design intent. Incorporating subjective human preferences and explicit design intent into alignment objectives remains a promising area for future work.
@@ 7 Conclusion
We demonstrated the adaption of alignment strategies from language modeling to preserve design intent in parametric CAD sketches. By using feedback from a constraint solver as a learning signal, we show the feasibility and value of alignment in parametric CAD tasks. In doing so, we pave the way for future AI-assisted design tools that incorporate design alignment.
@@ A.1 Background
A.2 Constraint State Definitions
The formal characterization of sketch constraint states follow Hoffman [17, 16]. Let a sketch be parameterized by a set of geometric parameters $P=\{p_{1},\dots,p_{n}\}$ and defined by a set of constraint equations $C=\{c_{1},\dots,c_{m}\}$. The Jacobian matrix $J_{C}=\frac{\partial C}{\partial P}$ captures the local dependency between parameters and constraints. A sketch is fully-constrained (FC) if $\text{rank}(J_{C})=n-r$, where $r$ represents the residual rigid-body degrees of freedom. It is under-constrained (UC) if $\text{rank}(J_{C})<n-r$, implying that at least one geometric parameter retains an unconstrained degree of freedom. It is over-constrained (OC) if $\text{rank}(J_{C})>n-r$, indicating redundant or inconsistent constraints, which may still yield a solvable configuration if the constraints are algebraically consistent.
A.3 Unstable Sketch Definition
To evaluate whether a sketch remains geometrically stable after constraint application, we introduce a metric that detects significant shifts in the sketch geometry. Specifically, we divide the sketch canvas into an $n\times n$ grid of spatial bins as shown in Figure A.2. A sketch is deemed unstable if any of its geometric entities move from their original bin to a different bin after constraint solving. This condition implies a meaningful deformation rather than minor numeric jitter. Such instability may indicate poorly conditioned constraint sets, where the solver resolves constraints by distorting the geometry. We apply this rule to all generated outputs and classify each sketch as either stable or unstable.
A.4 Fusion Sketch Representation
Figure A.2: Visualization of stable versus unstable sketches using a $4\times 4$ grid. Sketches with all points remaining in the same cell are considered stable (top), while those that move to a different cell are marked unstable (bottom).
The Fusion 360 Gallery sketch format [41] organizes sketch elements into a hierarchical, structured representation, wherein a sketch is defined by a set of parametric geometric primitives and a set of explicit constraints between those primitives. Each geometric primitive (line, arc, circle, point, etc.) is described by its intrinsic parameters (e.g., endpoint coordinates for a line, center and radius for a circle). Alongside the primitives, the sketch includes constraints (e.g., coincident points, perpendicular or parallel lines) that impose geometric relationships to be satisfied simultaneously. These constraints serve to preserve design intent: for instance, a coincidence constraint can lock the endpoint of a line onto a circle’s circumference, or an equal-length constraint can enforce that two segments remain the same length.
@@ A.5 Sketch Tokenization
A.6 SketchGraphs Dataset
In addition to the main paper that describes how the SketchGraphs dataset was filtered and converted, we provide additional details regarding the motivation and practical considerations of each step are provided here. The primary goal of these refinements is to produce a clean, representative subset of sketches and ensure each example aligns with standard engineering constraints.
A.6.1 Data Preprocessing
In Table A.1 we list out the supported constraint and dimension types in Onshape terminology that we included in the training data. Notably, we filter out less prevelant constraints (Symmetric, Normal, Pattern) and dimensions (CenterLine, Projected) to focus the learning task on the core geometric relationship types which form the backbone of sketch geometry. These filtered types represent higher-level constructs that can be equivalently modeled using more fundamental constraints and dimensions. For example, Symmetric constraints can be composed using a combination of Midpoint, Equal, and Collinear constraints. Similarly, Pattern constraints typically express repeated geometry with equal spacing, which can be reconstructed through a combination of Equal dimensions and manually replicated constraints. By removing these non-core constraints, we simplify the constraint vocabulary the model must learn while still covering the vast majority of design intent in sketches.
@@ A.6.1 Data Preprocessing
Constraint- and Dimension-Level Statistics
Type Frequency (%) Sketch Frequency (%)
Coincident 16.39% 31.13%
Horizontal 19.18% 64.20%
Vertical 11.16% 36.01%
Parallel 19.70% 42.26%
Perpendicular 13.56% 47.98%
Tangent 7.62% 13.47%
MidPoint 7.49% 20.79%
Equal 4.79% 13.20%
Concentric 0.11% 0.48%
Offset 43.15% 21.84%
Diameter 48.21% 25.37%
Radius 5.98% 4.57%
Linear 2.66% 1.93%
Angle 0.01% 0.01%
A.6.2 Processed Data Statistics
Table A.2 also summarizes the distribution of geometric constraints and dimensions in the dataset. The Type Frequency column reports the percentage of each constraint or dimension type relative to the total number of constraints or dimensions. The Sketch Frequency column shows the percentage of sketches in which at least one instance of the constraint or dimension appears.
We observe that commonly used geometric constraints such as Horizontal, Vertical, Parallel, and Coincident dominate the dataset, consistent with standard sketching practices in parametric CAD modeling. More specialized constraints like Concentric or Tangent appear less frequently, which aligns with their more limited use in practice.
For dimensions, Diameter and Offset are most frequent, as circular and offset features are prevalent in mechanical design. Radius, Linear, and especially Angle dimensions appear less often, consistent with their relatively specialized applications. These trends support the realism and representativeness of the dataset, suggesting it captures authentic usage patterns by human experts in professional CAD environments.
@@ Appendix B Architecture and Experiment Details
B.1 Experimental Setup
All experiments are conducted on an AWS P5.48xlarge instance. The instance is equipped with eight NVIDIA H100 GPUs (80 GB HBM3 memory per GPU), 192 vCPUs, and 2 TB of system memory.
A single epoch of RL training with the SketchGraphs dataset takes $\sim$3 days to train. This is primarily due to the frequent interactions with the CPU-based constraint solver and the fact that solve times can be highly varied. Roughly half of the training time is spent on GPU computation and half on detokenization and solver interaction. We expect custom optimizations could significantly reduce training time.
B.2 Constraint-Level Accuracy Evaluation
Evaluating constraint generation by direct constraint-level accuracy (i.e., exact matches between predicted and ground-truth constraints) is not meaningful for the constraint generation task. First, most sketches in the SketchGraphs dataset contain only a partial set of constraints defined by the original designer. Consequently, the ground-truth data does not necessarily represent the only valid or complete solution for fully constraining the sketch. Second, for a given geometric configuration, there often exist multiple valid constraint sets that can yield an equivalent, fully-constrained and stable sketch. For instance, the same geometry can be constrained either by a combination of horizontal and vertical constraints or by applying equivalent dimensional constraints, both of which are acceptable in practice. This makes exact constraint matching an unreliable indicator of functional correctness.
Instead, we evaluate generated constraint sets using functional metrics that better reflect real-world utility, as described in Section 3. These include whether the generated sketch is fully-constrained, stable, and solvable—metrics directly tied to the practical usability of the generated constraints in CAD workflows.
B.3 Vitruvion
@@
Our model generates constraints and dimensions as a structured token sequence, where each token represents a geometric primitive, constraint type, or dimensional relationship. This sequence-based representation allows the model to flexibly express a wide variety of parametric relationships. With a proper detokenization step, these sequences can be converted into standard constraint and dimension definitions supported by commercial CAD tools. As a result, the generated outputs are not limited to a specific platform and can be directly imported into widely used software such as Fusion, AutoCAD, Onshape, and SolidWorks, enabling seamless integration with existing design workflows.
B.4 Preference-based Optimization Algorithms
@@ B.5 RL algorithms
KL regularization $\beta$ - - 0.01
Policy clipping threshold $\epsilon$ - - 0.2